OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = -5 - 5 cos θ

OpenStudy (anonymous):

had u made the graph of that?

OpenStudy (anonymous):

No, I was told to do it without making a graph :/

OpenStudy (anonymous):

ok so to test for symmetric about the x axis u must replace (theta) by (-theta) and see if u get the same equation

OpenStudy (anonymous):

-5-5cos- θ ?

OpenStudy (anonymous):

i cant write theta so i will write (t) instead -5 - 5cos(t) replacing t by -t -5-5cos(-t) given that cos(-t) = cos(t) -5-5cos(-t) = -5-5cos(t) so the graph is symetric about the x axis

OpenStudy (anonymous):

What If i had r = 5 cos 3θ instead?

OpenStudy (anonymous):

it is the same have u tried to replace theta by -theta?

OpenStudy (anonymous):

5cos(3t) then replace 5cos(-3t) 5cos(-3t) = 5cos(3t) symetric about the x axis?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

here r the test for symetry in polar coordinates http://www.saylor.org/site/wp-content/uploads/2011/04/GRAPHING-IN-POLAR-COORDINATES.pdf